OpenStudy (anonymous):

write a function rule to describe each statement. 14. the cost in dollars of printing dollar bills when it costs 3.8cents to print a dollar. 15.the amount of money your earn mowing lawn at $15 per lawn. 16.the profit you make selling flowers at $1.50 each when each flower costs you $.80

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

if it costs 3.8 cents or 0.038 dollars to print a dollar, then how much will it cost to print x dollars?

OpenStudy (anonymous):

i dont know what im suppose to do

jimthompson5910 (jim_thompson5910):

1 dollar costs 0.038 dollars to print 2 dollars cost 2*0.038 = 0.076 dollars to print 3 dollars cost 3*0.038 = 0.114 dollars to print etc etc

jimthompson5910 (jim_thompson5910):

so x dollars cost _____ dollars to print

OpenStudy (anonymous):

so 15. X lawn time 15 is the amount money being made

jimthompson5910 (jim_thompson5910):

yep x times 15 or 15 times x short way to write it: 15x so the function is f(x) = 15x

OpenStudy (anonymous):

okay s for 16. f(x)=1.50-.80 ???

OpenStudy (anonymous):

for 14 how do we put it in to f(x)= ??

jimthompson5910 (jim_thompson5910):

14 is f(x) = 0.038x notice if you plug in x = 1, you get f(x) = 0.038 if you plug in x = 2, you get f(x) = 0.076 etc etc

OpenStudy (anonymous):

did i get number 16 right

jimthompson5910 (jim_thompson5910):

1.5 - 0.8 = ??

OpenStudy (anonymous):

.70

jimthompson5910 (jim_thompson5910):

so it's f(x) = 0.7x